OpenStudy (anonymous):

A rectangle is 6 km longer than it is wide. Find the dimensions of the rectangle if its area is 112 sq-km. Length= Width= Solving Polynomial Equations by Factoring

OpenStudy (sburchette):

Have you constructed your equations or assigned variables yet?

OpenStudy (anonymous):

no

OpenStudy (sburchette):

What if we let x = length. That would make the width = x-6 since the length is 6 km longer than the width. Since area is length * width, we have that x(x-6) = 112. Can you solve the equation from there?

OpenStudy (anonymous):

no

OpenStudy (sburchette):

We need to have the equation in factored form with zero on one side. If we distribute x on the left side, we get \[x^2-6x=112\] We can subtract 112 from both sides to obtain \[x^2-6x-112=0\]The left side factors into \[(x+8)(x-14)=0\] The solutions for x are -8 and 14. Since length cannot be negative, x must be 14. Since the width is 6 less than the length, the width is 8.